OpenStudy (anonymous):

Can someone explain to me how to substitute this equation into the Quad. formula? And PLEASE explain why you did what you did!: 4x^2-6x-10=0

OpenStudy (anonymous):

You could factor out a "2", but we won't do that for now. Your equation is in the form ax^2 + bx + c = 0 and you use\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]So, your a=4, b=-6, and c = -10. So, just substitute.

OpenStudy (anonymous):

But the +- part confuses me. =/

OpenStudy (anonymous):

np, the +- means you have (potentially) 2 different values for "x", depending on whether the expression under the radical is "0" or not. If the +- bothers you, you can write one equation with a + and another with a -. In time, you will not want to do that, because the notation +- is a time-saver that you will eventually become comfortable with.

OpenStudy (anonymous):

So, b = -6. That means -b = -(-6) = 6. Okay with that?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

This whole problem could be solved without the quadratic formula, but it would appear that the problem is posed so that you merely gain familiarity with the formula. Get to know this formula, as it is one of the very most important formulas in algebra. As for substitutions, can you substitute now?

OpenStudy (anonymous):

If a=4, b=-6, and c = -10 then we write that as 6+-14/8?

OpenStudy (anonymous):

If you mean (6 +- 14)/8, then yes, you are right! That 6 is also part of the numerator, and I think that's what you meant, but you have to put parentheses around it so it is not confused with 6 +- a number.

OpenStudy (anonymous):

Awesome. But do I write that as the final part of the equation?

OpenStudy (anonymous):

From (6 +- 14)/8, you would first factor out a 2 to make it look like (3 +- 7)/4, so you would have 5/2 or -1. Both numbers together make up your answer.

OpenStudy (anonymous):

as not yo confuse with sign, just use one operation at a time. conventionally, do addition first. afterwards, do subtraction. this thus results in two solution.

OpenStudy (anonymous):

Ahhh, I get it now. Again, thanks so much. :D

OpenStudy (anonymous):

You are quite welcome!